OpenStudy (anonymous):

I tried to do this one like the other ones together. 8+2i/2-6i

OpenStudy (anonymous):

I foiled (8+2i)(2+6i)=16+48i+4i+12i^2

OpenStudy (anonymous):

i got to 4+52i/40 and im stuck

OpenStudy (goformit100):



OpenStudy (anonymous):

ok so worked on it. 1/10+52i/40 =1/10+13/10

OpenStudy (anonymous):

\[\frac{ 1 }{ 10 }+\frac{ 13 }{ 10 }\]

OpenStudy (anonymous):

13/10i

jimthompson5910 (jim_thompson5910):

if you meant \[\Large \frac{ 1 }{ 10 }+\frac{ 13 }{ 10 }i\] (which I'm assuming you did), then you are correct

OpenStudy (anonymous):

wow im getting good! ty

jimthompson5910 (jim_thompson5910):

yes you are, very nice work